# Explanation Regularisation through the Lens of Attributions

### v1

@@ Abstract
Explanation regularisation (ER) has been introduced as a way to guide models to make their predictions in a manner more akin to humans, i.e., making their attributions ‘plausible’. This is achieved by introducing an auxiliary explanation loss, that measures how well the output of an input attribution technique for the model agrees with relevant human-annotated rationales. One positive outcome of using ER appears to be improved performance in out-of-domain (OOD) settings, presumably due to an increased reliance on ‘plausible’ tokens. However, previous work has under-explored the impact of the ER objective on model attributions, in particular when obtained with techniques other than the one used to train ER. In this work, we contribute a study of ER’s effectiveness at informing classification decisions on plausible tokens, and the relationship between increased plausibility and robustness to OOD conditions. Through a series of analyses, we find that the connection between ER and the ability of a classifier to rely on plausible features has been overstated and that a stronger reliance on plausible tokens does not seem to be the cause for any perceived OOD improvements. [^1] Source code available at https://github.com/PedroMLF/ER_through_the_lens_of_attributions.
@@
Pedro Ferreira 1  Wilker Aziz 1  Ivan Titov 1,2 1 University of Amsterdam  2 University of Edinburgh { p.m.ferreira, w.aziz }@uva.nlititov@inf.ed.ac.uk
@@ 1 Introduction
Figure 1: ER (top-left) minimises a classification and an explanation loss. The latter uses an input attribution technique to obtain a machine rationale for the prediction, and penalises differences between that and the human rationale. Guiding the model to attribute predictions to plausible tokens is expected to help the classifier after training (top-right), when no human rationale is available, and the classifier might be exposed to OOD inputs. Bottom: as any attribution technique is only a proxy to faithful input attributions, there can be a mismatch between predicting the human rationale (through the guided attribution technique) and effectively informing the classifier’s prediction by human rationale tokens.
@@
In explanation regularisation (ER; Ross et al., 2017; Ghaeini et al., 2019; Liu and Avci, 2019; Ismail et al., 2021; Joshi et al., 2022; Pruthi et al., 2022; Stacey et al., 2022, i.a.), a text classifier is encouraged to inform its predictions by tokens identified as part of a human rationale for the label. [^2] In this setting, a human rationale is the subset of input tokens that the annotator perceives as relevant to justify the label. This is achieved with an auxiliary ‘explanation loss’ that penalises differences between the output of a guided attribution technique (e.g., a relevance map based on top-layer attention) and the annotated human rationale (see Figure 1). Compared to their baseline counterparts trained without rationales, these ER models have been reported to improve classification in several out-of-domain (OOD) conditions, e.g., unseen test data (Joshi et al., 2022; Stacey et al., 2022; Madani and Minervini, 2023). This improved robustness is ascribed to increased reliance of ER models on plausible tokens (i.e., those in the human rationale for the label).
While ER is assumed to encourage classifiers to attribute their decisions more plausibly, and this increased plausibility is assumed to explain the observed OOD improvements, these have not been carefully verified in previous work. In particular, existing analyses have focused on the impact of explanation regularisation on guided attributions (i.e., those provided by the technique used in training), finding that ER models are better at aligning those with human rationales (Stacey et al., 2022; Madani and Minervini, 2023). However, through the explanation loss, guided attributions are explicitly supervised to resemble human rationales. Hence, we argue that this observation is an expected consequence of minimising the explanation loss, and not evidence that the classifier attributes its decisions more plausibly. Instead, we need to entertain the possibility of a faithfulness gap: the output of an attribution technique might not necessarily be aligned with the decision process of the model. Furthermore, ER can exacerbate this gap, making the guided attribution technique even less faithful.
In this work, we analyse (i) ER’s ability to make classifiers rely on plausible tokens; and (ii) the relationship between increased plausibility and robustness to OOD conditions. We start by studying in an in-domain setting how ER impacts input attributions using a range of post-hoc attribution techniques (§6.2). Here we find that, while ER can find models whose guided attributions effectively predict human rationales, there is little to no evidence of that being true under other (evaluation) post-hoc attribution techniques. This makes us doubt that ER substantially affects the model’s classification decision process. This finding persists even after making use of a potentially more faithful guided attribution technique (§6.3). Only by strongly constraining the model to achieve a low explanation loss in training, do we observe increased reliance on plausible features (§6.4), but that happens at the expense of OOD improvements. We follow this result by studying the trade-off between the classification and explanation losses (§6.5). Finally, we look into whether in-domain classification performance and plausibility of attributions available for model selection can be predictive of OOD classification improvements (§6.6). We observe that OOD improvements are difficult to predict, and that ER’s impact on the plausibility of input attributions of OOD examples is more limited than in-domain.
To summarise, we find that the connection between ER and the ability of a classifier to rely on plausible tokens has been overstated. While there is likely some effect connecting OOD classification improvements to classifiers that are better predictors of plausible tokens, the evidence we collect points to this effect not being a stronger reliance on plausible tokens.
@@ 2 Related Work
Human explanations, e.g., rationales, have been used to improve text classifiers (Hartmann and Sonntag, 2022; Hase and Bansal, 2022). Examples include select-predict ‘pipelines’, where an explanation output by a first module serves as input to a classifier (Camburu et al., 2018), and multi-task settings, where a classifier and an explanation predictor model are jointly trained (Chan et al., 2022). The latter can also be framed as a multi-objective optimization problem, as in Resck et al. (2024), a recent work on the trade-off between plausibility and in-domain classification performance in Pareto-optimal models.  In contrast, we are interested in explanation regularisation, where a model is trained such that its attributions, according to a differentiable attribution technique, match existing rationales (Ross et al., 2017; Ghaeini et al., 2019; Liu and Avci, 2019; Rieger et al., 2020; Mathew et al., 2021; Pruthi et al., 2022; Stacey et al., 2022, i.a.). Most work on ER focuses on in-domain data, however, there is evidence that ER is capable of improving robustness to OOD data, for both text Rieger et al. (2020); Joshi et al. (2022); Stacey et al. (2022), and image classification Chefer et al. (2022); Rao et al. (2023). We focus on this setup and further explore how ER impacts OOD attributions and performance.
@@
In vector-based techniques, model components are directly used to obtain input attributions. The simplest approach corresponds to using attention weights (Clark et al., 2019, i.a.), despite conflicting evidence on its usefulness for this purpose (Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019). Later works (Kobayashi et al., 2020, 2021; Ferrando et al., 2022; Modarressi et al., 2022) incorporate information about the magnitude of input vectors and the influence of other parts of the Transformer layer (Vaswani et al., 2017). These are examples of local techniques, where attributions are based on the dynamics of a single layer.
@@ 3 Explanation Regularisation
In ER (Joshi et al., 2022; Stacey et al., 2022; Pruthi et al., 2022, i.a.), we assume the availability of training data $\mathcal{D}=\{(x_{i},y_{i},e_{i})_{i=1}^{N}\}$, for each data point, $x_{i}$ is the input text, $y_{i}$ is the target label, and $e_{i}$ is a human rationale (specifically, a subset of tokens of the input regarded as a justification for the label). Human rationales can be obtained from annotators (Socher et al., 2013; Camburu et al., 2018; Rajani et al., 2019; Mathew et al., 2021) or through some form of heuristic to select relevant tokens (Liu and Avci, 2019; Rieger et al., 2020), and are only required during training.
@@
$\mathcal{L}_{\mathrm{cls}}(\theta|\mathcal{D})+\lambda~{}\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D}),$ (1)
@@
where $\mathcal{L}_{\mathrm{cls}}$ is the regular classification loss and $\mathcal{L}_{\mathrm{expl}}$ is the explanation loss, with a hyperparameter $\lambda\in\mathbb{R}_{\geq 0}$ controlling the contribution of the latter. The expectation is that $\mathcal{L}_{\mathrm{expl}}$ encourages the model to attribute classification decisions more similarly to human rationales, and this, in turn, leads to models that achieve better classification performance, particularly in out-of-domain conditions.
The first term in Eq. (1) is the cross-entropy loss
$\mathcal{L}_{\mathrm{cls}}(\theta|\mathcal{D})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\mathrm{ce}}(C_{\theta}(x),y)]~{},$ (2)
where $C_{\theta}$ is a trainable classifier. [^3] The typical $C_{\theta}$ employs a pre-trained text encoder and projects the top-layer representation of the [CLS] token to a probability distribution over all classes. The second term employs an attribution technique in order to obtain a relevance map $E(C_{\theta},x)$ for the classifier’s output. ER takes this as a proxy to faithful attributions, and penalises the classifier by this map’s deviation from the human rationale, as assessed by a function $\mathcal{S}$ such as mean absolute error, mean squared error, or KL divergence (Joshi et al., 2022):
$\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D})=\mathbb{E}_{(x,y,e)\sim\mathcal{D}}[\mathcal{S}(E(C_{\theta},x),e)].$ (3)
To avoid confusing the posthoc attribution technique that guides the model during training with other posthoc attribution techniques we may be interest in for analysis, we will refer to the former as the guided attribution technique and the latter as the evaluation attribution techniques. In addition to being differentiable, the guided attribution technique is required to be memory- and time-efficient, since it is used for training. Examples of such techniques used in ER include gradient-based (Ross et al., 2017; Ghaeini et al., 2019; Chefer et al., 2022), vector-based (Mathew et al., 2021; Pruthi et al., 2022; Stacey et al., 2022; Fernandes et al., 2022), and perturbation-based approaches (Ying et al., 2022). We follow Joshi et al. (2022) and Stacey et al. (2022) and use top-layer attention.
4 Method
We now outline two changes to the ER approach (§3) that provide the necessary level of empirical evidence to study our research questions.
4.1 Global Attributions as Supervision Signal
Intuitively, the guided attribution technique in ER controls the potential outcome of exploiting the additional supervision in human rationale data: if highly faithful to the classifier’s decision process, the model will be constrained to processing the input as if it were masked (i.e., to omit tokens the annotator deemed unessential), otherwise, the model will be able to process the input somewhat independently of what the guided attributions look like (i.e., of how human-rationale-looking they happen to be). A potential side effect of the former is to overly-constrain the classifier, forcing it to disregard portions of the input necessary for (machine) decision making. A potential side effect of the latter is to make no use of the additional supervision (e.g., the model might have enough capacity to minimise $\mathcal{L}_{\text{expl}}$ without any meaningful changes to the features available to the classification head). To study this, we employ a global attribution technique for guidance. For taking more of the model’s components into account when building a relevance map for the prediction, these techniques are potentially more faithful, or, as at least intuition suggests, they may lead to more substantial changes to how the input is processed at lower layers. We experiment with attention rollout (Abnar and Zuidema, 2020) for, on the one hand, it is a direct global counterpart to top-layer attention, and, on the other hand, it is much more memory- and time-efficient than vector-based alternatives (e.g., ALTI and DecompX). In Appendix A, we develop the intuition a bit more by comparing $\nabla_{\theta}\mathcal{L}_{\text{expl}}$ as a function of whether the guided attribution technique is local or global.
4.2 Constrained Optimization Approach
Suppose we are interested in analysing the best classifier we can get while meeting a constraint on how it attributes its decisions to input tokens. This is useful, for example, when studying the connection between reliance on plausible (human-rationale) tokens and OOD performance across different design choices (e.g., guided attribution technique). We can reinterpret the ER objective slightly and achieve that, by casting it as a constrained optimisation problem—minimise classification loss subject to a bound on explanation loss:
$\min_{\theta}$ $\quad\mathcal{L}_{\mathrm{cls}}(\theta|\mathcal{D})$ $\mathrm{s.t.}$ $\quad\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D})\leq b\;.$ (4)
For each choice of guided attribution technique, we can obtain a bound $b$ by interpreting $\mathcal{L}_{\text{expl}}$ as the sole objective of optimisation. To select the bound, we first independently minimise $\mathcal{L}_{\mathrm{expl}}$, setting $b$ to $\min_{\theta}~{}\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D})$. For Eq. (4), we use Lagrangian relaxation (Boyd and Vandenberghe, 2004). Implementation choices are detailed in Appendix F.
@@ 5 Experimental Setting
We use SST-2 (Socher et al., 2013) as training and in-domain data, following the heuristic algorithm proposed in Carton et al. (2020) to obtain instance-level rationales. For OOD we use: Amazon Reviews (Hou et al., 2024), IMBD (Maas et al., 2011), Movies (Zaidan and Eisner, 2008; DeYoung et al., 2020), and Yelp (Zhang et al., 2015). For more details on data refer to Appendix F.2.
Model
All experiments use HuggingFace Transformers’ (Wolf et al., 2020) BigBird-roberta-base model as the pre-trained contextual embedding encoder model (Zaheer et al., 2020). This choice follows previous works (Chan et al., 2022; Joshi et al., 2022; Madani and Minervini, 2023). For more training details refer to Appendix F.
@@
We use Mean Absolute Error (MAE) as the scoring function $\mathcal{S}$ in Equation 3.
Guided Attribution Technique.
We use top-layer attention scores (Att) as our local guided attribution technique, $E_{c}$, as explored before in the literature (Joshi et al., 2022; Stacey et al., 2022; Pruthi et al., 2022). As part of our analysis we also use attention rollout (AttR) as a global guided attribution technique. This technique recursively aggregates the local attention scores into a global representation. In both cases we use the scores for the [CLS] token, averaged across heads.
@@
All datasets with rationales are annotated at the instance level, with a binary vector with length equal to the number of tokens, where a value of one corresponds to a token that is highlighted as relevant. In order to obtain a target for the scoring function $\mathcal{S}$, the instance-level vector is normalized to sum to one by dividing each element by the total number of highlighted tokens in the instance (Joshi et al., 2022; Stacey et al., 2022).
ID OOD
Dev Test Movies Yelp IMDB Amazon-M-TV

### v3

@@ Abstract
Explanation regularisation (ER) has been introduced as a way to guide text classifiers to form their predictions relying on input tokens that humans consider plausible. This is achieved by introducing an auxiliary explanation loss that measures how well the output of an input attribution technique for the model agrees with human-annotated rationales. The guidance appears to benefit performance in out-of-domain (OOD) settings, presumably due to an increased reliance on ‘plausible’ tokens. However, previous work has under-explored the impact of guidance on that reliance, particularly when reliance is measured using attribution techniques different from those used to guide the model. In this work, we seek to close this gap, and also explore the relationship between reliance on plausible features and OOD performance. We find that the connection between ER and the ability of a classifier to rely on plausible features has been overstated and that a stronger reliance on plausible tokens does not seem to be the cause for OOD improvements.  [^1] Source code available at https://github.com/PedroMLF/ER_through_the_lens_of_attributions.
@@
Pedro Ferreira 1  Ivan Titov 1,2  Wilker Aziz 1 1 University of Amsterdam  2 University of Edinburgh { p.m.ferreira, w.aziz }@uva.nlititov@inf.ed.ac.uk
@@ 1 Introduction
Figure 1: ER (top-left) minimises a classification and an explanation loss. The latter uses an input attribution technique to obtain a machine rationale for the prediction, and penalises differences between that and the human rationale. Guiding the model to rely in its predictions on plausible tokens is expected to help the classifier at test time (top-right), when no human rationale is available. Bottom: even though the attribution technique used for guidance shows human-like rationales, the model may in fact rely on different (non-plausible) tokens, as other attribution techniques might reveal.
@@
In explanation regularisation (ER; Ross et al., 2017; Ghaeini et al., 2019; Liu and Avci, 2019; Ismail et al., 2021; Joshi et al., 2022; Pruthi et al., 2022; Stacey et al., 2022, i.a.), a text classifier is encouraged to inform its predictions by tokens included in a human rationale for the label. This is achieved with an auxiliary ‘explanation loss’ that penalises differences between the output of a guided attribution technique (e.g., a relevance map based on top-layer attention) and the annotated human rationale (see Figure 1). Compared to their counterparts trained without rationales, these ER models have been reported to improve classification, including for out-of-domain (OOD) inputs (Joshi et al., 2022; Stacey et al., 2022; Madani and Minervini, 2023), with this improved robustness being ascribed to increased reliance of ER models on plausible tokens (i.e., those in the human rationale for the label).
While ER is assumed to encourage classifiers to rely more on plausible tokens, this has not been carefully verified in previous work. In particular, prior analyses have focused on guided attributions—i.e., those explicitly supervised to resemble human rationales through the explanation loss—finding them to be better aligned with human rationales (Mathew et al., 2021; Stacey et al., 2022; Madani and Minervini, 2023, i.a.). Although impact on guided attributions is a necessary condition, this impact is not sufficient to demonstrate that the classifier relies more on plausible tokens. We argue that, to confirm an increased reliance, it is also necessary to examine the impact of guidance on attributions obtained through other non-guided attribution techniques (see Figure 1). If these are unaffected, it would suggest that the model is ‘hacking’ (Skalse et al., 2022) the explanation loss rather than genuinely increasing its reliance on plausible tokens.
In this work, we analyse (i) ER’s ability to make classifiers rely on plausible tokens; and (ii) the relationship between plausibility and robustness to OOD conditions. We start by studying ER with jointly optimised losses (§5.1). We find that – unlike the guided attribution techniques – there is little to no evidence that attributions by any other non-guided attributions are affected by ER, suggesting the loss ‘hacking’. In fact, we find that meaningful impact on input attributions is only achievable by ensuring a low explanation loss, for example via constrained optimisation (see §5.2). However, this is effective only for ‘global’ guided attribution techniques (which capture the information ‘flow’ across all or most components of the model), and, crucially, at the expense of classification performance.
Contributions.
We show that the connection between ER and the ability of a classifier to rely on plausible tokens has been overstated. In particular, we find that the choice of guided attribution technique is critical for this assumption: local attribution techniques, such as top-layer attention, can be ‘hacked’, minimising the explanation loss without much evidence of increased reliance on plausible tokens. Moreover, we find that OOD classification performance degrades for ER models that (over-)rely on plausible tokens, hinting at the need of finding better design choices for ER.
@@ 2 Related Work
Human explanations (e.g., token-level rationales, as in Figure 1) have been used to improve text classifiers (Hartmann and Sonntag, 2022; Hase and Bansal, 2022). Examples include select-predict ‘pipelines’, where an explanation output by a first module serves as input to a classifier (Camburu et al., 2018), and multi-task settings, where a classifier and a rationale extractor are jointly trained (Carton et al., 2022; Chan et al., 2022).   In contrast, we are interested in explanation regularisation, where, rather than having additional model components trained to perform rationale extraction, a model is trained to align attributions provided by a differentiable attribution technique with human rationales (Ross et al., 2017; Ghaeini et al., 2019; Liu and Avci, 2019; Rieger et al., 2020; Mathew et al., 2021; Pruthi et al., 2022; Stacey et al., 2022, i.a.). Most work on ER focuses on in-domain data, however, there is evidence that ER can improve robustness to OOD data, for text Rieger et al. (2020); Joshi et al. (2022); Stacey et al. (2022), and image classification Chefer et al. (2022); Rao et al. (2023). Other works identify difficulties in using explanations to learn better classifiers and propose methods to improve the compromise between classification performance and model explainability. These works differ from ours by: (i) not using human-annotated explanations (Plumb et al., 2020); or (ii) by departing significantly from the ER formulation, for example, by using a multi-task setup with a separate explainer (Carton et al., 2022) or by not employing a trainable explainer and instead incorporating human explanations through contrastive learning (Resck et al., 2024). We, instead, focus on ER and explore its impact on the plausibility of input attributions and how that, in turn, impacts classification performance both in- and out-of-domain.
@@
In vector-based techniques, model components are directly used to obtain input attributions. The simplest approach is to use attention weights (Clark et al., 2019, i.a.), despite conflicting evidence on its usefulness for this purpose (Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019; Pruthi et al., 2020). Later works (Kobayashi et al., 2020, 2021; Ferrando et al., 2022; Modarressi et al., 2022) incorporate information about the magnitude of input vectors and the influence of other components of the Transformer layer (Vaswani et al., 2017). These are examples of local techniques, where attributions are based on the dynamics of a single layer.
@@ 3 Explanation Regularisation
In ER (Joshi et al., 2022; Stacey et al., 2022; Pruthi et al., 2022, i.a.), we assume the availability of training data $\mathcal{D}=\{(x_{i},y_{i},e_{i})_{i=1}^{N}\}$, for each data point, $x_{i}$ is the input text, $y_{i}$ is the target label, and $e_{i}$ is a human rationale (specifically, a subset of tokens of the input regarded as a justification for the label). Human rationales can be obtained from annotators (Socher et al., 2013; Camburu et al., 2018; Rajani et al., 2019; Mathew et al., 2021) or through heuristics (Liu and Avci, 2019; Rieger et al., 2020), and are only required during training.
@@
$\mathcal{L}_{\mathrm{cls}}(\theta)+\lambda~{}\mathcal{L}_{\mathrm{expl}}(\theta),$ (1)
@@
where $\mathcal{L}_{\mathrm{cls}}$ is the regular classification loss and $\mathcal{L}_{\mathrm{expl}}$ is the explanation loss, with a hyperparameter $\lambda\in\mathbb{R}_{\geq 0}$ controlling the contribution of the latter. The expectation is that $\mathcal{L}_{\mathrm{expl}}$ leads the model to rely in its decisions on more plausible tokens.
The first term is the cross-entropy loss $\mathcal{L}_{\mathrm{cls}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\mathrm{ce}}(C_{\theta}(x),y)]$, where $C_{\theta}$ is a trainable classifier. The second term, $\mathcal{L}_{\mathrm{expl}}(\theta)=\mathbb{E}_{(x,y,e)\sim\mathcal{D}}[\mathcal{S}(E(C_{\theta},x),e)]$, employs an attribution technique in order to obtain a relevance map $E(C_{\theta},x)$ for the classifier’s output and penalises the classifier by this map’s deviation from the human rationale, as assessed by a function $\mathcal{S}$ such as mean absolute error or KL divergence (Joshi et al., 2022).   We refer to $E$ as the guided attribution technique. Besides being differentiable, $E$ is required to be memory- and time-efficient. Examples of techniques used in ER include gradient-based (Ross et al., 2017; Ghaeini et al., 2019; Chefer et al., 2022), vector-based (Mathew et al., 2021; Pruthi et al., 2022; Stacey et al., 2022; Fernandes et al., 2022), and perturbation-based approaches (Ying et al., 2022). We follow Joshi et al. (2022) and Stacey et al. (2022) and use top-layer attention and InputXGradient (Shrikumar et al., 2017). In addition, we also experiment with attention-rollout (Abnar and Zuidema, 2020).
@@ 4 Experimental Setting
We use SST-2 (Socher et al., 2013) as training and in-domain data, following the heuristic proposed by Carton et al. (2020) to obtain instance-level rationales. For OOD data with rationales we use Movies (Zaidan and Eisner, 2008; DeYoung et al., 2020), and annotate Yelp-50, a subset of Yelp, as described in Appendix G.1. For OOD without rationales annotations we use Amazon Reviews (‘Movies and TV’ split) (Hou et al., 2024), IMBD (Maas et al., 2011), and Yelp (Zhang et al., 2015). For more details on data refer to Appendix G.
Model.
All experiments use HuggingFace Transformers’ (Wolf et al., 2020) BigBird-roberta-base model as the pre-trained contextual embedding encoder model (Zaheer et al., 2020). This choice follows previous works (Chan et al., 2022; Joshi et al., 2022; Madani and Minervini, 2023). For more training details refer to Appendix H.
@@
We follow Joshi et al. (2022) and use mean absolute error for $\mathcal{S}$ in $\mathcal{L}_{\mathrm{expl}}(\theta)$.
@@
All datasets with rationales are annotated at the instance level with binary vectors, whose dimensionality equals the number of tokens. A value of 1 indicates relevant, highlighted tokens. In order to obtain a target for the scoring function $\mathcal{S}$, the instance-level vector is normalized to sum to one by dividing each element by the total number of highlighted tokens in the instance when using top-layer attention and attention-rollout (Joshi et al., 2022; Stacey et al., 2022). For InputXGradient we use the original binary values.
Attribution Techniques.
We make use of one local technique, namely, top-layer attention (Att), and 5 global techniques, namely, attention rollout (AttR; Abnar and Zuidema, 2020), InputXGradient (IxG; Shrikumar et al., 2017), ALTI Ferrando et al. (2022), and DecompX Modarressi et al. (2023) with (DX-C) and without (DX) the classification head.  [^2] For Att and AttR, we average attention weights across heads. For IxG we use the Captum package (Kokhlikyan et al., 2020). For DecompX and ALTI we adapt the original code. For guidance we use Att, AttR or IxG, as those are memory- and time-efficient enough to be used during training. The remainder are used only for analysis. All techniques are further described in Appendix A.
Plausibility Metrics.
To assess how well input attributions reproduce the human annotated rationales, we use three metrics introduced in Fomicheva et al. (2021), and described in Appendix D: AUC Score, Average Precision; and Recall@k.
Constrained Optimisation.
To study classifiers whose guided attributions are as plausible as they can be (through the lens of the explanation loss), we can reimagine ER as a constrained optimisation problem: minimise classification loss subject to a bound $b$ on the explanation loss,
$\min_{\theta}$ $\quad\mathcal{L}_{\mathrm{cls}}(\theta)$ $\mathrm{s.t.}$ $\quad\mathcal{L}_{\mathrm{expl}}(\theta)\leq b\;,$ (2)
with $b$ set to a value close to $\min_{\theta}~{}\mathcal{L}_{\mathrm{expl}}(\theta)$ – a minimiser of the explanation loss, taken in isolation. We approach this via Lagrangian relaxation (Boyd and Vandenberghe, 2004), with implementation choices detailed in Appendix H.
5 Results
ER Attr. F1 Guided Non-Guided
Joint5.1) Local $\varnothing$
Global $\varnothing$ $\varnothing$
Constr5.2) Local $\varnothing$
Global
$\times$
Table 1: Summary of the effect ($\checkmark$ positive,
$\times$
negative, or $\varnothing$ nil) of local or global guidance, across Joint and Constrained setups, in terms of OOD classification (F1) and impact on guided and non-guided attributions.
Table 1 presents a brief overview of our observations, which are discussed in detail in the following subsections. In §5.1, we find that despite positively impacting the OOD generalisation of the classification task, joint ER has limited impact on input attributions. In fact, only local guidance is able to better predict plausible rationales, and that effect is only observed for the guided technique itself. With global guidance, we find that the explanation loss is under-optimised. To understand what is happening, in §5.2 we reinterpret ER as constrained optimisation, where the explanation loss is constrained to be low. At this point, we find it to be possible to affect input attributions meaningfully, but only when guiding a global attribution technique and, crucially, at the expense of classification performance. In a nutshell, local guidance is being ‘hacked’: attributions are modified locally as to optimise the explanation loss without affecting the features used for classification (intuitively, the model ‘hides’ non-plausible computations where the local attribution technique cannot ‘see’, such as in lower layers). Global guidance, on the other hand, seems much harder to ‘hack’: optimising it well requires modifying computations performed in different layers across the model, more strongly restricting the features used for classification to rationale tokens. This, however, tends to worsen classification performance. Finally, we find in §5.3 that the disconnect between (over-)reliance on plausible tokens and OOD classification performance prevents us from systematically using features available for model selection to find models that perform best OOD.
5.1 Joint Optimisation
We start by studying the most common ER setup (§3), where the two losses are jointly optimised.
Joint ER improves OOD classification.
As observed in Table 2, joint ER improves OOD average classification performance. This is more noticeable for guidance with local attention (ER+Att), with guidance that uses global attribution techniques (ER+AttR and ER+IxG) showing smaller improvements. If we consider variance across runs, and visualise the distribution of results across seeds (Fig. 2), it is noticeable how the effect on the average improvement is exacerbated by runs that perform poorly. In fact, this seems to indicate that one of the merits of ER is to converge less often to models that generalise poorly to OOD conditions.
SST Dev SST Test Movies Yelp IMDB Amazon-M-TV